OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1. Show your work for full credit.

OpenStudy (anonymous):

will fan and medal!!

OpenStudy (jdoe0001):

notice the angle notice the the "x" value use SOH CAH TOA \(\bf sin(\theta)=\cfrac{y}{r} \\ \quad \\ % cosine cos(\theta)=\cfrac{x}{r} \\ \quad \\ % tangent tan(\theta)=\cfrac{y}{x}\implies tan\left( -\frac{3\pi }{4} \right)=\cfrac{y}{-1}\implies -1\cdot tan\left( -\frac{3\pi }{4} \right)=y\)